OpenStudy (anonymous):

Write the equation of a line that is perpendicular to the given line and that passes through the given point. 4x – 12y = 2; (10, –1) A. y = –3x + 29 B. y = -1/3x + 7 C. y=- 1/3x + 29 D. y = 3x + 29

OpenStudy (kat_s123):

I gotchu! Soo do you know what the slope of a line that is perpendicular to another line will look like?

OpenStudy (kat_s123):

mmkay, well, the first thing you'd want to do is put the given equation into y intercept form. Which is y = mx +b.So In order to do that we want to get y alone.

OpenStudy (kat_s123):

so you'd subtract 4x from both sides of the equation, which leaves you with -12y = -4x +2. Next thing you'd need to do is divide both sides by -12. That would give you y = 4/12x -2/12. So now all you need to do is simplify it. y = 1/3x - 1/6 So now we now what the slope is - 1/3.

OpenStudy (kat_s123):

Soo the rule with perpendicular lines is that the line that it is perpendicular to has to have the opposite reciprocate. So for this specific problem the perpendicular slope would be -3. And only one answer has -3 as its slope, so there's your answer!